# Review of deep learning models for crypto price prediction: implementation and evaluation

### v1

@@ 5 Discussion
In this section, we provide a discussion based on the results, taking into consideration the model architecture as well as the characteristics of the data. In summary, we evaluated the predictive performance of all models and presented the results through Figures 14 to 21. We also provide a ranking of the performance accuracy for each type of model in Tables 12 and 13.
@@
We first review the results of the first experiment that investigated the model performance with COVID-19 data. Our results show that LSTM, BD-LSTM and ED-LSTM provide outstanding predictive performance across four different cryptocurrencies and two different approaches for selecting model variables. The CNN, Conv-LSTM, and Transformer models show good performance only under particular conditions. We also note that in all the cryptocurrencies, the Univariate model outperformed the Multivariate model, but in some cases (Bitcoin and Dogecoin) the accuracy was close when comparing both strategies. We found that the models with high forecast accuracy were mostly accompanied by narrower confidence intervals. On the contrary, the higher RMSE values usually resulted in lower robustness of the model. We found that models with better prediction accuracy (lower RMSE) provide more robust performance accuracy, given different model weight initializations in independent experimental runs. The accuracy of performance generally declines as the prediction horizons increase, which is natural for multistep ahead problems (Figure 14b). The prediction is derived from the current values, and the information gap expands as the prediction step increases. This is because our task is defined as a direct approach for forecasting multisteps, rather than an iterated prediction strategy. We observe the changes in the prediction horizon of each model and find that CNN and Conv-LSTM are significantly worse than other models. The forecast accuracy of these two models frequently declines more rapidly than that of other models, and occasionally even show volatility (Figure 18b), across Step 1 to 5. We found that the CNN-related models using convolution operation provided lower accuracy than other models in predicting cryptocurrency price. Later we will analyse the cause of this issue. Among the predictions for the four currencies in Experiment 1 (Table 10), Dogecoin has the worst prediction effect. The RMSE values of the model are significantly higher than those of the other three cryptocurrencies. We believe this is due to the particularity of the Dogecoin price, which leads to large prediction errors. The first 70% of the data fluctuates smoothly, while the last 30% of the data fluctuates violent.
@@
In the second experiment, we evaluate the accuracy deep learning model predicting cryptocurrencies. We utilize the two models that predicted the best in previous experiments to do forecasting with the new dataset. The new dataset includes all close prices since COVID-19 to April 2024. It has been discovered that the close-price forecasts for Bitcoin, Dogecoin, and Ethereum have all shown enhancement. Also, as the prediction horizon rises, the prediction accuracy of the model deteriorates more slowly. We claim that there are two causes contributing to the decrease in the accuracy of forecasting litecoin. The first is that because our evaluation criterion for the model is the overall performance of the model, we did not choose Conv-LSTM, which had the best performance in predicting Litecoin in previous experiment. Maybe Conv-LSTM is more suitable for predicting Litecoin. The second reason is that Litecoin had a violent price ups and downs before COVID-19. Due to our design of the experiment, the data during this period was not included in the training set.
@@
Next, we aim to investigate what might be contributing to the lower accuracy of multivariate models compared to univariate models in prediction. In our analysis (Fig or Table xxx, we notice that the price of cryptocurrencies is extremely unstable and is greatly influenced by several variables outside or inside of the market [10]. Simply inserting some additional factors will not only be ineffective in assisting the model to accurately forecast outcomes, but it may also mislead the model into acquiring knowledge of irrelevant data features.
Our analysis of volatility (Figure 10 and 11) shows the high degree of volatility exhibited by cryptocurrencies throughout the COVID-19 pandemic. Through a comparative analysis of the outcomes obtained from first and second experiment, we observed that the use of high volatility data as the training set provides better prediction accuracy for the model. The robustness and scalability of the model are also improved.
@@
Next, we investigate the factors contributing to the advantages and disadvantages of each model. Conventional time series models and machine learning techniques are inadequate for addressing issues such as timing dependency and gradient explosion. As we used MLP and ARIMA models to predict Bitcoin, the prediction performance was not as good as the deep learning model. And according to [139], there is long-term memory in the cryptocurrency market. The LSTM network, a deep learning model initially designed to address long-term memory issues, distinguishes itself in this regard. The memory gate in the LSTM network can better capture information in time series with long-term dependencies. The prediction performance of CNN is worse than that of LSTM, which is what we expected. Because the convolutional layer in CNN is better at capturing local patterns and features, it has better prediction results for sequences where there is obvious spatial correlation between data points. According to past research, CNN seems to be more effective in handling image recognition problems. Next, we analyze the differences between the four models with LSTM layers (LSTM, ED-SLTM, BD-LSTM and Conv-LSTM). The ED-LSTM model has been created for language modeling problems, particularly for sequence to sequence modeling in language translation. In this model, an encoder LSTM is used to transform a source sequence into a fixed-length vector, while a decoder LSTM is employed to convert the vector representation back into a variable-length target sequence [105]. In our study, the encoder function maps an input time series to a vector of fixed length. Subsequently, the decoder LSTM function translates the vector representation to several prediction horizons. Despite the differences in the application, the fundamental objective of mapping inputs to outputs remains unchanged. As a result, ED-LSTM models have proven to be highly efficient for multi-step ahead prediction. BD-LSTMs utilize two LSTM models to capture both forward and backward information about the sequence at each time step [102]. While these models have been shown effective for language modeling, our findings indicate that they’re useful in tracking both present and future states for time series modeling. In our experiments, BD-LSTM and ED-LSTM provided stable and outstanding prediction performance. Although Conv-LSTM uses convolutional layers as input with LSTM memory cells in the hidden layer, it differs from the conventional LSTM models since the memory cells from different hidden layers update only in the time domain and are mutually independent [140]. Therefore, information at the top layer in time $t-1$ will be ignored by the bottom layer at time $t$. In cryptocurrency price prediction, the time information is crucial. This also explains why the prediction accuracy of Conv-LSTM in our experiments is often low and unstable at high prediction horizons. We also employed the Transformer provided unsatisfactory results which may be attributed to the limited training data, as Transformer models are often better suited for handling large amounts of data.
@@ 6 Conclusions
In this study, we provide a rigorous evaluation of novel deep learning models for cryptocurrency price forecasting. We compared prominent deep learning models using univariate and multivariate strategies. The results show that the Bidirecional-LSTM provides the highest accuracy in predicting cryptocurrency prices. We also provided a comparison with baseline models such as multilayer perceptron and ARIMA and found that deep learning models generally outperform them. We also found that multivariate models provided less prediction efficiency than univariate models; however, it has scope for improvement given the availability of higher correlated time series data as features. In terms of the effect of COVID-19, we found that close-price volatility for cryptocurrency is quite apparent. Our experimental results show that utilising a training data set with high volatility enhances the precision of our predictions.
In future work, it would be worthwhile to improve the multivariate model. It is advisable to utilise more dependable factors to enhance forecasts, maybe employing techniques like causal inference to find these variables. We can also use this study framework to switch the goal into predictions of other financial indicators such as volatility of cryptocurrency. Further applications to other specific issues could also be viable, such as predicting energy use and extreme weather forecasting.

### v2

@@ 5 Discussion
We first reviewed the results of the first experiment that investigated the model performance without COVID-19 data. Our results show that Conv-LSTM with Multivariate* strategy provided outstanding predictive performance across four different cryptocurrencies. We also note that in all the cryptocurrencies, the Univariate model was poorer than the Multivariate* strategy. The Univariate Strategy was generally better in accuracy than the Multivariae strategy for Experiment 1 (Tables 10 and 11). We found that the models with high forecast accuracy(lower RMSE) were mostly accompanied by narrower confidence intervals (eg. Figure 22). On the contrary, the higher RMSE values usually resulted in lower robustness of the model, ie. higher confidence interval. The accuracy of performance generally declines as the prediction horizons increase, which is natural for multistep ahead problems (Figure 14b). The prediction is derived from the current values, and the information gap expands as the prediction step increases. This is because our task is defined as a direct approach for forecasting multisteps, rather than an iterated prediction strategy. We found that the CNN-related models using convolution operation provided better accuracy than other models in predicting cryptocurrency price. Later we will analyse the cause of this issue. Among the predictions for the four currencies in Experiment 1 (Table 11), Dogecoin has the worst prediction effect. The RMSE values of the model are significantly higher than those of the other three cryptocurrencies. We believe this is due to the particularity of the Dogecoin price, which leads to large prediction errors. The first 70% of the data fluctuates smoothly, while the last 30% of the data fluctuates highly.
@@
In Experiment 2, we evaluated the two strategies (Univariate and Multivariate*) for the dataset featuring COVID-19 and beyond. It has been discovered that the close-price forecasts for Bitcoin, Dogecoin, and Ethereum have all shown improvements. Also, as the prediction horizon rises, the prediction accuracy of the model deteriorates more slowly. We claim that there are two causes contributing to the decrease in the accuracy of forecasting Litecoin. The first is that because our evaluation criterion for the model is the overall performance of the model, we did not choose Conv-LSTM, which had the best performance in predicting Litecoin in the previous experiment. Maybe Conv-LSTM is more suitable for predicting Litecoin. The second reason is that Litecoin was highly volatile before COVID-19. Due to our design of the experiments, the data during this period was not included in the training set.
@@
Next, we aim to investigate what might be contributing to the lower accuracy of multivariate models compared to univariate models in prediction. In our analysis (Figure 10 and 11), we noticed that the price of cryptocurrencies is extremely unstable and is greatly influenced by several variables outside or inside of the market [10]. Simply inserting some additional factors will not only be ineffective in assisting the model to accurately forecast outcomes, but it may also mislead the model into acquiring knowledge of irrelevant data features. Our analysis of volatility (Figure 10 and 11) shows the high degree of volatility exhibited by cryptocurrencies throughout the COVID-19 pandemic. In both experiments (Tables 10, 11 and 12), we observed that the use of a highly correlated crypto as a feature in the Multivariate* strategy provided better prediction accuracy for the model.
@@
Next, we investigate the factors contributing to the advantages and disadvantages of each model. Conventional time series models and machine learning techniques are inadequate for addressing issues such as timing dependency and gradient explosion. As we used MLP and ARIMA models to predict Bitcoin, the prediction performance was not as good as the deep learning model. Jiang et al. [139] highlighted that there is a long-term memory in the cryptocurrency market. The memory gate in the LSTM network can better capture information in time series with long-term dependencies and the convolutional layer in CNN is better at capturing local patterns and features for multivariate datasets. Hence, the combination of CNN with LSTM (Conv-LSTM), provided the best results. Next, we analyze the differences between the four models with LSTM layers (LSTM, ED-SLTM, BD-LSTM and Conv-LSTM). The ED-LSTM model has been created for language modeling problems, particularly for sequence to sequence modeling in language translation. In this model, an encoder LSTM is used to transform a source sequence into a fixed-length vector, while a decoder LSTM is employed to convert the vector representation back into a variable-length target sequence [105]. In our study, the encoder function maps an input time series to a vector of fixed length. Subsequently, the decoder LSTM function translates the vector representation to several prediction horizons. Despite the differences in the application, the fundamental objective of mapping inputs to outputs remains unchanged. As a result, ED-LSTM models have proven to be highly efficient for multi-step ahead prediction. BD-LSTMs utilize two LSTM models to capture both forward and backward information about the sequence at each time step [102]. Although these models (BD-LSTM and LD-LSTM) have been shown effective for language modelling, our findings indicate that they’re useful in both univariate and multivariate strategies and provided competitive performance with Conv-LSTM. Although Conv-LSTM uses convolutional layers as input with LSTM memory cells in the hidden layer, it differs from the conventional LSTM models since the memory cells from different hidden layers update only in the time domain and are mutually independent [140]. Therefore, information at the top layer in time $t-1$ will be ignored by the bottom layer at time $t$. In cryptocurrency price prediction, the time information is crucial. This also explains why the prediction accuracy of Conv-LSTM in our experiments is often low and unstable at high prediction horizons. The Transformer model that has enhanced attention mechanism provided unsatisfactory results in Experiment 2 (lowest rank in Table 14) and Experiment 1 (lowest rank in Table 13 for Univariate and Multivariate strategies. However, the Transformer model was number 2 rank in the case of Multivariate* for Experiment 1 which suggests that it works better with data that has less vitality (pre COVID-19). The issue with poor results may be attributed to the limited training data, as Transformer models are often better suited for handling large amounts of data.
Our findings indicate that combining the closing prices of the highest-correlated cryptocurrencies into the multivariate model leads to a significant enhancement in the model’s accuracy in forecasting. The Conv-LSTM model provides the most accurate predictions when employing this strategy for forecasting cryptocurrencies for both experiments. Under the Multivariate* approach, the scalability and accuracy of models becomes better. This means that there is minimal difference in the accuracy of predicting 1 day and predicting 5 days. Generally, we found that Experiment 2 has lower model accuracy than Experiment 1, since the latter features data during COVID-19 which features phases of high votality.
The major limitation of the framework deals with a lack of uncertainty quantification which can be addressed by Bayesian deep learning, which includes MCMC sampling and variational inference-based deep learning models . Furthermore, data augmentation methods can be explored to exploit segments of highly correlated crypto and stick prices, from different historical periods. We featured the prediction of daily price, but the model can be extended for regular hourly price prediction and also be extended to include extreme value forecast strategies, such as quantile regression . Furthermore, enhanced Bayesian data imputation methodology can extrapolate missing data in cases better when highly correlated data sources have such issues. Finally, we only looked at unimodal data. Due to the complexity of the cryptocurrency market, further extension can be done using a multimodal data approach that considers text data from news media and social media using large language models .
@@ 6 Conclusions
In this study, we provide a rigorous evaluation of novel deep learning models for cryptocurrency price forecasting. We compared prominent deep learning models using univariate and multivariate strategies. The results show that the convolutional LSTM with multivariate strategy provides the highest accuracy in predicting cryptocurrency prices for two major experimental setup featuring data, period to and during COVID-19 pandemic. Additionally, our findings indicate that the multivariate strategy featuring a highly correlated cryptocurrency data enhanced prediction accuracy significantly, in comparison to the univariate strategy. In terms of the effect of COVID-19, we found that close-price volatility for cryptocurrency is quite apparent, which brings further challenges to the respective models. Our experimental results show that utilising a training data set with high volatility weakens the precision of our predictions.
In future work, it would be worthwhile to improve the multivariate model further using more correlated factors to enhance forecasts by employing techniques such as causal inference to find more correlated data in the multivariate strategy. We can also use a framework for predictions of other financial indicators, such as the volatility of cryptocurrency prices and market stocks. Further applications to other domains, such as predicting energy use and extreme weather forecasting, can also be done.